OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (-2, 1). Which equation defines circle P?

OpenStudy (anonymous):

(x + 2)2 + (y – 1)2 = 8 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64

OpenStudy (anonymous):

use equation of a circle

OpenStudy (anonymous):

hmm can you tell me the equation?

OpenStudy (anonymous):

(x + 2)2 + (y – 1)2 = 64

OpenStudy (anonymous):